Muriatic Acid, an industrial grade of concentrated Hydrochloric Acid is used to clean masonry and cement. Given that the concentration is 11.7 M, how many milliliters of muriatic acid will contain 9.66 g of HCl?
1
Expert's answer
2020-09-24T08:13:00-0400

Let volume of solution is V ml having concentration 11.7 M


So , moles of solute =( V * 11.7 / 1000 ) mol.


V * 11.7 / 1000 = 9.66 / 36.5


V = 22.620 ml . Answer .